# Diffusion Meets DAgger: Supercharging Eye-in-hand Imitation Learning

### v1

@@ Abstract
A common failure mode for policies trained with imitation is compounding execution errors at test time. When the learned policy encounters states that were not present in the expert demonstrations, the policy fails, leading to degenerate behavior. The Dataset Aggregation, or DAgger approach to this problem simply collects more data to cover these failure states. However, in practice, this is often prohibitively expensive. In this work, we propose Diffusion Meets DAgger (DMD), a method to reap the benefits of DAgger without the cost for eye-in-hand imitation learning problems. Instead of collecting new samples to cover out-of-distribution states, DMD uses recent advances in diffusion models to create these samples. This leads to robust performance from few demonstrations. In experiments conducted for non-prehensile pushing on a Franka Research 3, we show that DMD can achieve a success rate of 80% with as few as 8 expert demonstrations, where behavior cloning reaches only 20%. DMD also outperform competing NeRF-based augmentation schemes by 50%.
@@ I Introduction
Imitation learning is an effective way to train robots. However, even when testing policies in environments similar to those used for training, imitation learning suffers from the well-known Compounding Execution Errors problem: small errors made by the learned policy lead the robot to out-of-distribution states causing the learner to make even bigger errors (compounding errors) [56].
One solution to this problem is via manual collection of expert-labeled data on states visited by the learner, a strategy popularly known as Dataset Aggregation or DAgger [56]. In theory, this solution works. However, putting it into practice is challenging from a practical stand point: driving a robot that doesn’t obey your commands is difficult. Many alternatives have been proposed, e.g. [35, 36], but they all require collecting more expert data. In this paper, we pursue an alternate paradigm: automatically generating observations and action labels for out-of-distribution states. By replacing data collection with data creation, we improve the sample efficiency of imitation learning. In fact, data creation was Pomerleau’s original solution to this problem [51]. We revisit, improve, and automate his solution from 30 years ago using modern data-driven image generation methods. We do this for imitation learning problems in context of eye-in-hand setups (i.e. setups where images come from a camera mounted on the robot hand) that are increasingly becoming more popular [64, 84, 77].
@@
In more detail: given a behavior cloning trajectory, $\tau$ collected from an expert operating the system, we design generative models to synthesize off-distribution states and corresponding action labels. For example, for the object pushing application in Figure 1, this would correspond to generating off-center views of the object from the good expert trajectory $\tau$. Specifically, we learn a function $f(I_{t},\Delta p)$ that synthesizes the observation $\tilde{I}_{t}$ at a small perturbation $\Delta p$ to the trajectory at time step $t$. Knowledge of this $\Delta p$ lets us compute the ground truth action label $\tilde{a}_{t}$ for this out-of-distribution observation. We augment the expert demonstration $\tau$ with multiple off-distribution samples $(\tilde{I}_{t},\tilde{a}_{t})$ along the trajectory. Access to out-of-distribution states and action labels for how to recover from them improves the performance of the task policy.
In this paper, we develop and evaluate this idea in context of eye-in-hand manipulation policies in the real world. While the idea is intuitive, a number of nuances have to be dealt with carefully to get it to work. As we have access to the entire trajectory, the natural choice is to adopt NeRFs [48, 85] to realize the function $f$. While NeRFs work very well for view synthesis in computer vision, we found that they don’t work well for the synthesis task at hand because of the inevitable deformations in the scene as the gripper manipulates the objects. We thus switched to using diffusion models for this synthesis task. This led to good image synthesis even in situations where the scene deformed over the course of manipulation. Computing action labels for these samples presented yet another challenge (Figure 6). The labels should align with making progress towards the goal. We thus investigated different schemes for sampling $\Delta p$’s and computing action labels $\tilde{a}_{t}$.
This paper presents experiments that evaluate the aforementioned design choices to develop a data creation framework to supercharge eye-in-hand imitation learning. Experiments on a physical platform demonstrate the effectiveness of our proposed framework over baseline data augmentation techniques. On a non-prehensile pushing on a Franka Research 3, we show that DMD can achieve a success rate of 80% with as few as 8 expert demonstrations, where behavior cloning reaches only 20%. We also outperform a recently proposed augmentation scheme based on NeRFs [85] by 50%.
@@ II-A Imitation Learning
Behavior cloning (BC), or training models to mimic expert behavior, has been a popular strategy to train robots over the last many decades [50, 51, 58] and has seen a recent resurgence in interest [28, 86]. Ross et al. [56] theoretically and empirically demonstrate the compounding execution error problem (small errors in the learner’s predictions steer the agent into out-of-distribution states causing the learner to make even bigger errors) and propose a Dataset Aggregation (DAgger) strategy to mitigate it. Over the years, researchers have sought to improve the basic BC and DAgger recipe in several ways [25, 4]. [64, 77, 17, 84, 62] devise ways to ease and scale-up expert data collection, while [83] design ways to collect demonstrations in virtual reality. [35, 36] devise ways to simplify DAgger data collection. [14, 61] model the multiple modes in expert demonstrations, while [49] employ a non-parameteric approach. Researchers have also shown the effectiveness of pre-trained representations for imitation learning [49, 74]. [85, 12, 46, 81] employ image and view synthesis models to synthetically augment data. Complementary to our work, [12, 46, 81] focus on generalization across objects via semantic data augmentation by adding, deleting and editing objects in the scene.
[85, 31, 16, 32] pursue synthetic data augmentation to address the compounding execution error problem. [31, 16] conduct these augmentation in low dimensional state spaces. Ke et al. [32] leverage local continuity in the environment dynamics to generate corrective labels for low-dimensional state spaces and LiDAR observations in simulation. Closest to our work, Zhou et al. [85]’s SPARTN model train NeRF’s on demonstration data to synthesize out-of-distribution views. NeRF assumes a static scene. Thus, SPARTN can’t reliably synthesize images when the scene undergoes deformation upon interaction of the robot with the environment. Our use of diffusion models to synthesize images gets around this issue and experimental comparison to SPARTN demonstrate the effectiveness of our design choice.
@@ II-B Image and View Synthesis Models
Recent years have witnessed large progress in image generation [18, 24, 54, 63, 65]. This has led to a number of applications: text-conditioned image generation [54, 5, 20, 57, 52, 82], image and video inpainting [44, 9, 47], image to image translation [27] and generation of novel views for objects and scenes from one or a few images [72, 69, 80, 42, 34]. While diffusion models have proven effective at image synthesis tasks, Neural Radiance Fields (NeRFs) [48] excel at view interpolation tasks. Given multiple (20 – 100) images of a static scene, vanilla NeRFs can effectively interpolate among the views to generate photo-realistic renders. Researchers have pursued NeRF extensions that enable view synthesis from few images [79, 19, 7, 75] and even model deforming scenes [66, 8]. In general, NeRFs are more effective at interpolation problems, while diffusion models excel at problems involving extrapolation (e.g. synthesizing content not seen at all) or where exactly modeling changes in 3D is hard (e.g. deforming scenes). As our work involves speculating how the scene would look if the gripper were at a different location, we adopt a diffusion based methods. We show comparison against a NeRF based method as well.
@@ II-C Diffusion Models and Neural Fields for Robot Learning
Effective generative models and neural field based representations have been used in robotics in other ways than for data augmentation. Researchers have used diffusion models for representing policies [37, 1, 38, 13, 71], generating goals or subgoals [30, 6, 10], offline reinforcement learning [3, 71, 43], planning [40, 29, 73, 41], behaviorally diverse policy generation [23], and predicting affordances [76]. Additionally, they have been used to generate samples with high reward [26] and skill aquisition from task or play data [11]. Neural fields have been used to represent scenes for manipulation [39, 70] and navigation [2].
@@ III Approach
Figure 2: DMD System Overview: Our system operates in three stages. a) A diffusion model is trained, using task and play data, that can synthesize novel views of a scene relative to a given image. b) This diffusion model is used to generate an augmenting dataset that contains off-trajectory views ($\tilde{I_{2}^{1}}$, $\tilde{I_{2}^{2}}$) from expert demonstrations. Labels for these views ( cyan arrows) are constructed such that off-trajectory views will still converge towards task success (right). Images with a green border are from trajectories in the original task dataset. Purple-outlined images are diffusion-generated augmenting samples. c) The original task data and augmenting dataset are combined for policy learning.
@@ III-A Overview
Given task data $\mathbb{D}$, imitation learning learns a task policy $\pi$. The task data takes the form of a collection of trajectories $\tau_{i}$, where each trajectory in turn is a collection of image action pairs, $(I_{t},a_{t})$. $\pi$ is trained using supervised learning to regress action $a_{t}$ from images $I_{t}$ over the task data $\mathbb{D}$. For the non-prehensile manipulation of objects using an eye-in-hand cameras considered in this paper, images $I_{t}$ correspond to views from the eye-in-hand camera as input, and the actions $a_{t}$ are the relative end-effector poses between consecutive time steps.
When trained with few demonstrations, $\pi$ exhibits poor generalization performance. To address poor generalization, our approach generates an augmented dataset $\mathbb{\tilde{D}}$ and trains the policy jointly on $\mathbb{\tilde{D}}\cup\mathbb{D}$. Augmented samples are specifically generated to be out-of-distribution from $\mathbb{D}$ and thus aid the policy’s generalization capabilities. $\mathbb{\tilde{D}}$ is generated through a conditional diffusion model $f$ that generates a $\Delta p$ perturbed view $\tilde{I}$ of an image $I_{t}$ from a given trajectory $\tau$: $\tilde{I}=f(I_{t},\Delta p)$. We use the action labels in the trajectory $\tau$ to compute the action label $\tilde{a}$ for this perturbed view. We describe the design of the conditional diffusion model in Section III-B and detail the augmenting image sampling and action labels computation procedure in Section III-C.
@@ III-B Diffusion Model for Synthesizing $\Delta p$ Perturbed Views
The function $f(I,\Delta p)$ is realized using a conditional diffusion model that is pretrained on large Internet-scale data. Specifically, we adopt the recent work from Yu et al. [80] and finetune it on data from $\mathbb{D}$. The model produces image $I_{b}$ by conditioning on a reference image $I_{a}$ taken by camera $a$ and a transformation matrix ${}_{a}T_{b}$ that maps points in camera $b$’s frame to camera $a$’s frame. By representing the desired $\Delta p$ as the transformation between two cameras (${}_{a}T_{b}=\Delta p)$, and using images from $\mathbf{D}$ as the reference images ($I_{a}\sim\mathbf{D}$), the learned diffusion model can generate the desired perturbed views for augmenting policy learning.
@@ III-B1 Model Architecture
A diffusion model can be thought of as an iterative denoiser: given a noisy image $x_{t}$, diffusion timestep $t$, it is trained to predict the noise $\epsilon$ added to the image.
$\mathcal{L}=\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2}$
@@ III-B2 Model Training
Figure 4: Visualization of perturbed images generated by diffusion model for augmenting policy learning. In each row, we show $\tilde{I_{t}}$ generated from $I_{t}$ through different camera translations.
Training the model requires access to triples: $(I_{a},I_{b},{}_{a}T_{b})$. We process data from $\mathbb{D}$ to generate these triples. Let each trajectory consist of images $I_{1},\dots,I_{N}$. We use the structure from motion [21, 68] implementation from the COLMAP library [59, 60] to extract poses for the images in the trajectory $\tau$. This associates each image $I_{t}$ in the trajectory with the (arbitrary) world frame ${}_{t}T_{w}$. We can then compute relative pose between arbitrary images $I_{a}$ and $I_{b}$ from a trajectory via ${}_{a}T_{b}={}_{a}T_{w}{}_{w}T_{b}$. We sample random pair of images from the trajectory to produce triples $(I_{a},I_{b},{}_{a}T_{b})$ that are used to train the model $\epsilon_{\theta}$. COLMAP also outputs the camera intrinsics necessary for training the model. Since COLMAP reconstruction is only correct up to a scale factor, we normalize translations by the largest displacement between adjacent frames per trajectory.
As $\mathbb{D}$ contains much less data than is typically required to train a diffusion models from scratch, we finetune the model from Yu et al. [80]. We find that finetuning with as few as 24 trajectories leads to realistic novel view synthesis for our task as shown in Figure 4. We only finetune the diffusion model weights and use the pre-trained VQ-GAN codebooks. We experiment with different datasets for finetuning this diffusion model: task data from $\mathbb{D}$, play data [45, 78] collected in our setup, and their combinations.
@@ III-C1 Sampling Out-of-distribution Images
We generate out of distribution views for images in the task dataset $\mathbf{D}$. $\Delta p$’s are sampled as follows. For an image $I_{t}$ from trajectory $\tau$, we sample vectors in a random direction, with magnitude drawn uniformly from $[0.2s,s]$, where $s$ is the largest displacement between adjacent frames in $\tau$. $\Delta p$ simply corresponds to translation along this randomly chosen vector. We constrain these samples to not be too aligned with the actual action.

### v2

@@ Abstract
A common failure mode for policies trained with imitation is compounding execution errors at test time. When the learned policy encounters states that are not present in the expert demonstrations, the policy fails, leading to degenerate behavior. The Dataset Aggregation, or DAgger approach to this problem simply collects more data to cover these failure states. However, in practice, this is often prohibitively expensive. In this work, we propose Diffusion Meets DAgger (DMD), a method that reaps the benefits of DAgger but without the cost, for eye-in-hand imitation learning problems. Instead of collecting new samples to cover out-of-distribution states, DMD uses recent advances in diffusion models to synthesize these samples. This leads to robust performance from few demonstrations. We compare DMD against behavior cloning baseline across four tasks: pushing, stacking, pouring, and hanging a shirt. In pushing, DMD achieves 80% success rate with as few as 8 expert demonstrations, where naive behavior cloning reaches only 20%. In stacking, DMD succeeds on average 92% of the time across 5 cups, versus 40% for BC. When pouring coffee beans, DMD transfers to another cup successfully 80% of the time. Finally, DMD attains 90% success rate for hanging shirt on a clothing rack.
@@ I Introduction
Imitation learning is an effective way to train robots for new tasks. However, even when testing policies in environments similar to those used for training, imitation learning suffers from the well-known Compounding Execution Errors problem: small errors made by the policy lead to out-of-distribution states, causing the robot to make even bigger errors (compounding errors) [56].
One solution to this problem is via manual collection of expert-labeled data on states visited by the learner, a strategy popularly known as Dataset Aggregation or DAgger [56]. However, DAgger [56] is challenging to put into practice: it requires an expert operator to supervise the robot during execution and guide it to recover from failures. Many alternatives have been proposed, e.g. [35, 36], but they all require collecting more expert data. In this paper, we pursue an alternate paradigm: automatically generating observations and action labels for out-of-distribution states. By replacing data collection with data creation, we improve the sample efficiency of imitation learning. In fact, data creation was Pomerleau’s original solution to this problem [52]. We revisit, improve, and automate his solution from 30 years ago using modern data-driven image generation methods.
@@
We target imitation learning problems in the context of eye-in-hand setups (i.e. setups where images come from a camera mounted on the robot hand), which are becoming increasingly more popular [64, 85, 78]. Given a behavior cloning trajectory $\tau$ collected from an expert, we design generative models to synthesize off-distribution states and compute corresponding action labels. For example, in the object pushing application shown in Figure 1, this corresponds to generating off-center views of the object from the good expert trajectory $\tau$. Specifically, we learn a function $f(I_{t},\Delta p)$ that synthesizes the observation $\tilde{I}_{t}$ at a small perturbation $\Delta p$ to the trajectory at time step $t$. $\Delta p$ lets us compute the ground truth action label $\tilde{a}_{t}$ for this out-of-distribution observation. We augment the expert demonstration $\tau$ with multiple off-distribution samples $(\tilde{I}_{t},\tilde{a}_{t})$ along the trajectory.
As we have access to the entire trajectory, one option is to realize $f(I_{t},\Delta p)$ using NeRFs [48, 86]. While NeRFs work very well for view synthesis in computer vision, we find they are not suitable for the synthesis task at hand because of the inevitable deformations in the scene as the gripper manipulates the objects. We thus switch to using diffusion models, and this change leads to high-quality image synthesis even when the scene deforms during manipulation. Computing action labels for these samples present yet another challenge (Figure 5). The labels should align with progress towards the goal. We thus investigate different schemes for sampling $\Delta p$’s and computing action labels $\tilde{a}_{t}$.
We present experiments that evaluate the aforementioned design choices in developing a data creation framework to supercharge eye-in-hand imitation learning. The framework is tested in four settings: non-prehensile pushing, stacking, pouring, and hanging a shirt. Across all tasks, we see a sizeable improvement over vanilla behavior cloning, demonstrating the effectiveness of our framework Diffusion Meets DAgger (DMD). For the task of non-prehensile pushing, DMD achieves 80% success rate given only 8 expert demonstrations, while behavior cloning (BC) reaches only 20%. It also outperforms a NeRF-based augmentation method [86] by 50%. For the task of stacking cups on a box, DMD achieves a success rate of 93% on average across three training cups, and 85% success for two unseen cups. In comparison, BC only reaches 36% and 45% success rates, respectively. For the task of pouring coffee beans into a cup, DMD succeeds 80% of the time, while BC falls short at 30%. Finally, DMD achieves a success rate of 90% for the hanging shirt task, improving upon the 20% success rate for BC.
@@ II-A Imitation Learning
Behavior cloning (BC), or training models to mimic expert behavior, has been a popular strategy to train robots over the last many decades [51, 52, 58], and has witnessed renewed interest in recent times [29, 88]. [56] theoretically and empirically demonstrate the compounding execution error problem (small errors in the learner’s predictions steer the agent into out-of-distribution states causing the learner to make even bigger errors) and propose a Dataset Aggregation (DAgger) strategy to mitigate it.
Over the years, researchers have sought to improve the basic BC and DAgger recipe in several ways [27, 4]. [64, 78, 19, 85, 62] devise ways to ease and scale-up expert data collection, while [84] design ways to collect demonstrations in virtual reality. [35, 36] devise ways to simplify DAgger data collection. [14, 61] model the multiple modes in expert demonstrations, while [49] employ a non-parametric approach. Researchers have also shown the effectiveness of pre-trained representations for imitation learning [49, 75]. [86, 13, 46, 82] employ image and view synthesis models to synthetically augment data. Complementary to our work, [11, 46, 82] focus on generalization across objects via semantic data augmentation by adding, deleting and editing objects in the scene.
[86, 32, 18, 50] pursue synthetic data augmentation to address the compounding execution error problem.  [32, 18, 50] conduct these augmentation in low dimensional state spaces. [50] develops a principled way to sample these augmentations. Closest to our work, Zhou et al. [86]’s SPARTN model train NeRF’s on demonstration data to synthesize out-of-distribution views. NeRF assumes a static scene. Thus, SPARTN can’t reliably synthesize images when the scene undergoes deformation upon interaction of the robot with the environment. Our use of diffusion models to synthesize images gets around this issue and experimental comparison to SPARTN demonstrate the effectiveness of our design choice.
@@ II-B Image and View Synthesis Models
Recent years have witnessed large progress in image generation [20, 26, 54, 63, 65]. This has led to a number of applications: text-conditioned image generation [54, 5, 22, 57, 53, 83], image and video inpainting [44, 10, 47], image to image translation [28] and generation of novel views for objects and scenes from one or a few images [73, 70, 81, 42, 34]. While diffusion models have proven effective at image synthesis tasks, Neural Radiance Fields (NeRFs) [48] excel at view interpolation tasks. Given multiple (20 – 100) images of a static scene, vanilla NeRFs can effectively interpolate among the views to generate photo-realistic renders. Researchers have pursued NeRF extensions that enable view synthesis from few images [80, 21, 7, 76] and even model deforming scenes [67, 9]. In general, NeRFs are more effective at interpolation problems, while diffusion models excel at problems involving extrapolation (e.g. synthesizing content not seen at all) or where exactly modeling changes in 3D is hard (e.g. deforming scenes). As our work involves speculating how the scene would like if the gripper were at a different location, we adopt a diffusion based methods. We show comparison against a NeRF based method as well.
@@ II-C Diffusion Models and Neural Fields for Robot Learning
Effective generative models and neural field based representations have been used in robotics in other ways than for data augmentation. Researchers have used diffusion models for representing policies [37, 1, 38, 14, 72], generating goals or subgoals [31, 6, 11], offline reinforcement learning [3, 72, 43], planning [40, 30, 74, 41], behaviorally diverse policy generation [25], predicting affordances [77], and skill acquisition from task or play data [12]. Neural fields have been used to represent scenes for manipulation [39, 71] and navigation [2].
Figure 2: DMD System Overview: Our system operates in three stages. a) A diffusion model is trained, using task and play data, to synthesize novel views relative to a given image. b) This diffusion model is used to generate an augmenting dataset that contains off-trajectory views ($\tilde{I_{2}^{1}}$, $\tilde{I_{2}^{2}}$) from expert demonstrations. Labels for these views (cyan arrows) are constructed such that off-trajectory views will still converge towards task success (right). Images with a green border are from trajectories in the original task dataset. Purple-outlined images are diffusion-generated augmenting samples. c) The original task data and augmenting dataset are combined for policy learning.
@@ III-A Overview
Given task data $\mathbb{D}$, imitation learning learns a task policy $\pi$. The task data comprises a set of trajectories $\tau_{i}$, each consisting of a sequence of image action pairs, $(I_{t},a_{t})$. Policy $\pi$ is trained using supervised learning to regress action $a_{t}$ from images $I_{t}$. In this paper, we consider manipulation of objects using an eye-in-hand camera, where input images $I_{t}$ correspond to views from a wrist camera, and the actions $a_{t}$ are the relative end-effector poses between consecutive time steps.
When trained with few demonstrations, $\pi$ exhibits poor online performance upon encountering states not covered by the expert demonstrations. To address this issue, as shown in Figure 2, our approach generates an augmented dataset $\mathbb{\tilde{D}}$ and trains the policy jointly on $\mathbb{\tilde{D}}\cup\mathbb{D}$. Augmented samples are specifically generated to be out-of-distribution from $\mathbb{D}$, thus helping the policy to generalize. $\mathbb{\tilde{D}}$ is generated through a conditional diffusion model $f$ that synthesizes a $\Delta p$-perturbed view $\tilde{I}_{t}$ of an image $I_{t}$ from a given trajectory $\tau$: $\tilde{I}_{t}=f(I_{t},\Delta p)$. We use action labels in the trajectory $\tau$ to compute the action label $\tilde{a}_{t}$ for this perturbed view. The design of the conditional diffusion model is described in Section III-B, and the procedure for sampling augmenting images and computing action labels is detailed in Section III-C.
@@ III-B Diffusion Model for Synthesizing $\Delta p$ Perturbed Views
The function $f(I_{t},\Delta p)$ is realized using a conditional diffusion model that is pretrained on large Internet-scale data. Specifically, we adopt the recent work from Yu et al. [81] and finetune it on data from $\mathbb{D}$. As shown in Figure 2(a), the model produces image $I_{b}$ by conditioning on a reference image $I_{a}$ taken by camera $a$ and a transformation matrix ${}_{a}T_{b}$ that maps points in camera $b$’s frame to camera $a$’s frame. By representing the desired $\Delta p$ as the transformation between two cameras (${}_{a}T_{b}=\Delta p)$, and using images from $\mathbf{D}$ as the reference images ($I_{a}\sim\mathbf{D}$), the learned model can generate the desired perturbed views for augmenting policy learning.
@@ III-B1 Model Architecture
A diffusion model is an iterative denoiser: given a noisy image $x_{t}$, diffusion timestep $t$, it is trained to predict the noise $\epsilon$ added to the image:
$\mathcal{L}=\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2}.$
@@ III-B2 Model Training
Training the model requires access to triples: $(I_{a},I_{b},{}_{a}T_{b})$. We process data from $\mathbb{D}$ to generate these triples. Let each trajectory consist of images $I_{1},\dots,I_{N}$. We use structure from motion (SfM) algorithms [23, 69, 59, 60, 8, 15] to extract poses for the images in the trajectory $\tau$. This associates each image $I_{t}$ in the trajectory with the (arbitrary) world frame ${}_{t}T_{w}$. We can then compute relative pose between arbitrary images $I_{a}$ and $I_{b}$ from a trajectory via ${}_{a}T_{b}={}_{a}T_{w}{}_{w}T_{b}$. We sample random pair of images from the trajectory to produce triples $(I_{a},I_{b},{}_{a}T_{b})$ that are used to train the model $\epsilon_{\theta}$.
Figure 4: Training Examples from the Diffusion Model and Computed Labels: We visualize generated examples, $\tilde{I}$, used to train our policies along with the computed action label (the arrow is a projection of the 3D action into the 2D image plane: the arrow pointing up means move the gripper forward, pointing to the right means move it right). The first row shows augmenting samples for the pushing task, while the second row shows those for the stacking task.
As $\mathbb{D}$ contains much fewer data than is typically required to train a diffusion models from scratch, we finetune the model from Yu et al. [81]. Finetuning with around 50 trajectories leads to realistic novel view synthesis for our tasks as shown in Figure 7. We only finetune the diffusion model weights. The pre-trained VQ-GAN codebooks is kept fixed.
The diffusion model can be finetuned on task data $\mathbb{D}$, task-agnostic play data [45, 79], or even combinations of task and play data. Play data in our setups involves moving the grabber within the workspace and randomly interacting with the task objects without emphasizing task completion. Access to random interactions in play data may make it easier for the diffusion model to synthesize diverse perturbations. We experiment with these different data sources for training the diffusion model (Section IV-A4 and Section IV-C2).
@@ III-C1 Sampling Out-of-distribution Images
With the noise estimation model $\epsilon_{\theta}$ trained, the image generation function $f(I_{t},\Delta p)$ is realized by simply letting $I_{a}=I_{t}$, letting ${}_{a}T_{b}$ represent the desired $\Delta p$, and sampling from the conditional diffusion model using well-established sampling strategies [66, 26, 54].
We use $f(I_{t},\Delta p)$ to generate out-of-distribution images with a very simple strategy for sampling $\Delta p$. For an image $I_{t}$ from trajectory $\tau$, we sample vectors in a random direction, with magnitudes drawn uniformly from a pre-defined range for each task. If the SfM reconstruction recovers the real-world scale, this range is set to $[2cm,4cm]$; otherwise, it is $[0.2s,s]$, where $s$ is the largest displacement between adjacent frames in $\tau$. $\Delta p$ simply corresponds to moving along this randomly chosen vector.
Figure 5: The Overshooting Problem: When the generated image $\tilde{I}_{t}$ exceeds $I_{t+1}$, the inferred action for $\tilde{I}_{t}$ may direct the agent away from task success. We refer to this as the overshooting problem. At time step $t+1$, the view $I_{t+1}$ has moved to the lower right of $I_{t}$. However, the synthesized sample $\tilde{I}_{t}$ has moved even further to the right than $I_{t+1}$, but not beyond $I_{t+3}$. Blue arrow represents action label for $\tilde{I}_{t}$ computed using $I_{t+1}$ as the target; green arrow represents action label computed using $I_{t+3}$ as the target. Since $\tilde{I}_{t}$ has overshot $I_{t+1}$, an action taken with $I_{t+1}$ as the next intended target moves backward, away from the apple, and this labeling is not desirable. Computing the action with respect to a farther image, say $I_{t+3}$, does not have this issue.